madmax2008:

Solve for x: 3x − 24 = 81

swinn:

3x - 24 = 81 3x = 81 + 24 3x = 105 divide both sides by 3 3x/3 = 105/3 so then x=

madmax2008:

thank you swinn 😁😁

swinn:

np